# The distribution on permutations induced by a random parking function

### v3

@@ Abstract.
parking function, random permutation
@@ 1. Introduction and Statement of Results
(1.1) $P_{n}^{\text{park}}(\sigma)=P^{\mathcal{P}_{n}}\left(T_{n}^{-1}(\sigma)\right).$
@@
For example, if $\sigma=379218645\in S_{9}$, then $l_{n,i}(\sigma)=1$, for $i\in\{1,4,5,7,8\}$, $l_{n,i}(\sigma)=2$, for $i\in\{2,9\}$ and $l_{n,i}(\sigma)=3$, for $i\in\{3,6\}$. Thus, $L_{9}(\sigma)=1^{5}2^{2}3^{3}=36$. At certain points in the paper, it will be convenient to use the following variant of $l_{n,i}$. Define
@@
We have the following proposition.
Proposition 1.
@@
The following corollary of Proposition 1 is immediate, where the asymptotic behavior follows from Stirling’s formula.
@@
We will prove the following weak convergence result for $L_{n}$.
@@
The following corollary follows immediately from Proposition 1 and Theorem 1.
@@
And the following corollary follows immediately from Corollary 1, Proposition 1 and Stirling’s formula, $n!\sim n^{n}e^{-n}\sqrt{2\pi n}$.
@@
Comparing (1.8) and (1.9), we see that for all but an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$, the $P_{n}^{\text{park}}$-probability of a permutation in $S_{n}$ is approximately $\frac{2^{n}}{(n+1)^{n-1}}$, but the “average” $P_{n}^{\text{park}}$-probability of a permutation in $S_{n}$ is exponentially larger, namely asymptotic to $\frac{e^{n}}{\sqrt{2\pi}n^{n+\frac{1}{2}}}$. By Stirling’s formula, it also follows that there is a set $A_{n}\subset S_{n}$ with $P_{n}(A_{n})\to 1$ for which $P_{n}^{\text{park}}(A_{n})$ is around $(\frac{2}{e})^{n}$. There is an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$ each of whose elements has super-exponentially larger $P_{n}^{\text{park}}$-probability than the above average probability, and an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$ for which the $P_{n}^{\text{park}}$-probability is exponentially smaller than the average probability. In particular, we have the following corollary.
@@
Let $\sigma^{-1}$ denote the inverse permutation of $\sigma$. So $\sigma^{-1}_{k}=j$ if and only if $\sigma_{j}=k$. From the definition of a parking function, it is obvious that
@@
The distribution $\{e^{-j}\frac{j^{j-1}}{j!}\}_{j=1}^{\infty}$ is called the Borel distribution. It is not obvious that it is a distribution, that is, that it sums to 1. For more on this, see [1, p.135]. It follows that (1.10) also holds with $P^{\mathcal{P}_{n}}$ and $\pi_{1}$ replaced by $P_{n}^{\text{park}}$ and $\sigma^{-1}_{1}$. It would be nice to obtain similar type asymptotics for $P_{n}^{\text{park}}(\sigma^{-1}_{j}=k)$, for general $j,k$. We pursue this direction when $j$ is near $n$, that is for the last cars to park.
@@
Immediate from (1.11). ∎
@@
Immediate from (1.11). ∎
@@
The proof of Proposition 1 is given in section 2 and the proof of Theorem 1 is given in section 3. The proof of Theorem 2 is given in section 4, the proof of Corollary 9 is given in section 5 and the proof of Theorem 3 is given in section 6. The proof of Proposition 2 is given in section 7.
2. Proof of Proposition 1
Recall the definitions of $l_{n,i}$ and $\tilde{l}_{n,i}$ in (1.2) and (1.3). For the proof of the proposition, it will be convenient to work with $\tilde{l}_{n,i}$. Recall from (1.4) that
@@ 2. Proof of Proposition 1
The proposition will follow if we show that for each $\sigma\in S_{n}$, there are $L_{n}(\sigma)$ different parking functions $\pi\in\mathcal{P}_{n}$ such that $T_{n}(\pi)=\sigma$, where $T_{n}$ is as in the paragraph containing equation (1.1). Before giving a formal proof of the proposition, we illustrate the proof with a concrete example, from which the general result should be clear. Consider the permutation $\sigma=379218645\in S_{9}$. We look for those $\pi\in\mathcal{P}_{9}$ that satisfy $T_{9}(\pi)=\sigma$. From the definition of the parking process and from the definition of $T_{n}$, we need $\pi_{1}=5$ in order to have $\sigma_{5}=1$, $\pi_{2}=4$ in order to have $\sigma_{4}=2$, $\pi_{3}=1$ in order to have $\sigma_{1}=3$ and $\pi_{4}=8$ in order to have $\sigma_{8}=4$. In order to have $\sigma_{9}=5$, we can either have $\pi_{5}=9$, in which case car number 5 parks in its preferred space 9, or alternatively, $\pi_{5}=8$, in which case car number 5 attempts to park in its preferred space 8 but fails, and then moves on to space 9 and parks. Then we need $\pi_{6}=7$ in order to have $\sigma_{7}=6$. Then similar to the explanation regarding $\pi_{5}$, we need $\pi_{7}$ to be either 1 or 2 in order to have $\sigma_{2}=7$. In order to have $\sigma_{6}=8$, we can have either $\pi_{8}=6$, in which car number 8 parks directly in its preferred space 6, or alternatively $\pi_{8}=5$, in which case car number 8 tries and fails to park in space number 5 and then parks in space number 6, or alternatively, $\pi_{8}=4$, it which case car number 8 tries and fails to park in space number 4 and then also in space number 5, before finally parking in space number 6. Similarly, we need $\pi_{9}$ to be equal to 1,2 or 3 in order to have $\sigma_{9}=3$. Thus, there are $1\times 1\times 1\times 1\times 2\times 1\times 2\times 3\times 3=\prod_{i=1}^{9}\tilde{l}_{9,i}(\sigma)=L_{9}(\sigma)$ different parking functions $\pi\in\mathcal{P}_{9}$ that yield $T_{9}(\pi)=\sigma$.
@@
To give a formal proof for the general case, fix $\sigma\in S_{n}$. In order to have $T_{n}(\pi)=\sigma$, first we need $\pi_{1}=\sigma^{-1}_{1}$. Thus there is just one choice for $\pi_{1}$, and note that $\tilde{l}_{n,1}(\sigma)=1$. Now let $k\in[n-1]$ and assume that we have chosen $\pi_{1},\cdots\pi_{k}$ in such a way that car number $i$ has parked in space $\sigma^{-1}_{i}$, for $i\in[k]$. We now want car number $k+1$ to park in space $\sigma^{-1}_{k+1}$. By construction, this space is vacant at this point, and so are the $\tilde{l}_{n,k}(\sigma)-1$ spaces immediately to the left of this space. However the space $\tilde{l}_{n,k}$ spaces to the left of this space is not vacant (or possibly this space doesn’t exist–it would be the zeroth space). Thus, by the parking process, car number $k+1$ will park in space $\sigma^{-1}_{k+1}$ if and only if $\pi_{k+1}$ is equal to one of the $\tilde{l}_{n,k+1}(\sigma)$ numbers $\sigma^{-1}_{k+1},\sigma^{-1}_{k+1}-1,\cdots,\sigma^{-1}_{k+1}-\tilde{l}_{n,k+1}(\sigma)+1$. This shows that there are $L_{n}(\sigma)=\prod_{i=1}^{n}\tilde{l}_{n,i}(\sigma)$ different parking functions $\pi$ satisfying $T_{n}(\pi)=\sigma$. $\square$
@@ 3. Proof of Theorem 1
Using (3.12), a standard straightforward calculation gives
@@ 4. Proof of Theorem 2
We continue counting in this fashion until we have chosen cars for the $k_{j}-k_{j-1}-1$ spaces between spaces $k_{j-1}$ and $k_{j}$. Recall that $k_{0}$ has been defined to be 0. Now we need for all of the $n-j-\sum_{l=1}^{j}(k_{i}-k_{i-1}-1)=n-k_{j}$ cars that remain from among the first $n-j$ cars to park in the spaces $k_{j}+1,\cdots,n$. In order for this to occur, but for no one of the first $n-j$ cars to park in the $k_{j}$th space, these remaining members of $\{\pi_{i}\}_{l=1}^{n-j}$ must constitute a parking function of length $n-k_{j}$, but shifted forward by $k_{j}$ spaces. This gives $(n-k_{j}+1)^{n-k_{j}-1}$ different choices.
@@
Now we count how many ways we can choice the $\{\pi_{l}\}_{l=n-j+1}^{n}$ in order that $\sigma_{n-j+l}=i_{l},\ l=1,\cdots,j$, or equivalently, in order that car $n-j+l$ park in space $i_{l},\ l=1,\cdots,j$. This calculation uses the same reasoning that was used in the proof of Proposition 1. After the first $n-j$ cars have parked appropriately as above, the spaces $\{k_{l}\}_{l=1}^{j}=\{i_{l}\}_{l=1}^{j}$ are still vacant. From the definition of the parking process and from the definition of $\tilde{l}_{n,n-j+1}(\sigma)$, in order that car $n-j+1$ park in space $i_{1}$, $\pi_{n-j+1}$ must take on one of the values $i_{1}.i_{1}-1,\cdots i_{1}-\tilde{l}_{n,n-j+1}(\sigma)+1$. Then in order that car $n-j+2$ park in space $i_{2}$, $\pi_{n-j+2}$ must take on one of the values $i_{2},i_{2}-1,\cdots i_{2}-\tilde{l}_{n,n-j+2}(\sigma)+1$. Continuing like this, we conclude that there are $\prod_{l=1}^{j}\tilde{l}_{n,n-j+l}(\sigma)$ choices. Using this with (4.1), and noting that there are $(n+1)^{n-1}$ different parking functions of length $n$, we obtain (1.11). $\square$

### v4

@@ Abstract.
parking function, random permutation, Borel distribution
@@ 1. Introduction and Statement of Results
(1.1) $P_{n}^{\text{park}}(\sigma)=P^{\mathcal{P}_{n}}\left(T_{n}^{-1}(\{\sigma\})\right).$
@@
For example, if $\sigma=379218645\in S_{9}$, then $l_{n,i}(\sigma)=1$, for $i\in\{1,4,5,7,8\}$, $l_{n,i}(\sigma)=2$, for $i\in\{2,9\}$ and $l_{n,i}(\sigma)=3$, for $i\in\{3,6\}$. Thus, $L_{9}(\sigma)=1^{5}2^{2}3^{2}=36$. At certain points in the paper, it will be convenient to use the following variant of $l_{n,i}$. Define
@@
We have the following theorem.
Theorem 1.
@@
The following corollary of Theorem 1 is immediate, where the asymptotic behavior follows from Stirling’s formula, $n!\sim n^{n}e^{-n}\sqrt{2\pi n}$.
@@
Although $E_{n}L_{n}$ is around $e^{n}$, with high $P_{n}$-probability $L_{n}(\sigma)$ is near $2^{n}$. We will prove the following weak convergence result for $L_{n}$.
@@
The following corollary follows immediately from Theorems 1 and 2.
@@
And the following corollary follows immediately from Theorem 1, Corollary 1 and Stirling’s formula.
@@
Comparing (1.8) and (1.9), we see that for all but an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$, the $P_{n}^{\text{park}}$-probability of a permutation in $S_{n}$ is approximately $\frac{2^{n}}{(n+1)^{n-1}}$, but the “average” $P_{n}^{\text{park}}$-probability of a permutation in $S_{n}$ is exponentially larger, namely asymptotic to $\frac{e^{n}}{\sqrt{2\pi}n^{n+\frac{1}{2}}}$. By Stirling’s formula, it also follows that there is a set $A_{n}\subset S_{n}$ with $P_{n}(A_{n})\to 1$ for which $P_{n}^{\text{park}}(A_{n})$ is around $(\frac{2}{e})^{n}$. There is an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$ each of whose elements has super-exponentially larger $P_{n}^{\text{park}}$-probability than the average probability, and an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$ each of whose elements has exponentially smaller $P_{n}^{\text{park}}$-probability than the average probability. In particular, we have the following corollary.
@@
Let $\sigma^{-1}$ denote the inverse permutation of $\sigma$. So $\sigma^{-1}_{k}=j$ if and only if $\sigma_{j}=k$. In car parking language,$\sigma^{-1}_{k}=j$ means that car number $k$ parked in space number $j$. From the definition of a parking function, it is obvious that
@@
The distribution $\{e^{-j}\frac{j^{j-1}}{j!}\}_{j=1}^{\infty}$ is called the Borel distribution. It is not obvious that it is a distribution, that is, that it sums to 1. For more on this, see [1, p.135]. It follows that (1.10) also holds with $P^{\mathcal{P}_{n}}$ and $\pi_{1}$ replaced respectively by $P_{n}^{\text{park}}$ and $\sigma^{-1}_{1}$. It would be nice to obtain asymptotic results for $P_{n}^{\text{park}}(\sigma^{-1}_{j_{n}}=k_{n})$, for general $j_{n},k_{n}$. We pursue this direction when $j_{n}$ is near $n$, that is for the last cars to park.
@@
Immediate from (1.11), noting that if $\sigma^{-1}_{n}=k$, then $\tilde{l}_{n,n}(\sigma)=k$. ∎
@@
Immediate from (1.11), noting that if $\sigma^{-1}_{n-1}=l$ and $\sigma^{-1}_{n}=m$, then
$\begin{cases}\tilde{l}_{n,n-1}(\sigma)=l,\ \tilde{l}_{n,n}(\sigma)=m,\ \text{if}\ l<m;\\ \tilde{l}_{n,n-1}(\sigma)=l-m,\ \tilde{l}_{n,n}(\sigma)=m,\ \text{if}\ l>m.\end{cases}$
@@
The proof of Theorem 1 is given in section 2 and the proof of Theorem 2 is given in section 3. The proof of Theorem 3 is given in section 4, the proof of Corollary 9 is given in section 5 and the proof of Theorem 4 is given in section 6. The proof of Proposition 1 is given in section 7.
2. Proof of Theorem 1
Recall the definitions of $l_{n,i}$ and $\tilde{l}_{n,i}$ in (1.2) and (1.3). For the proof of the theorem, it will be convenient to work with $\tilde{l}_{n,i}$. Recall from (1.4) that
@@ 2. Proof of Theorem 1
The theorem will follow if we show that for each $\sigma\in S_{n}$, there are $L_{n}(\sigma)$ different parking functions $\pi\in\mathcal{P}_{n}$ such that $T_{n}(\pi)=\sigma$, where $T_{n}$ is as in the paragraph containing equation (1.1). Before giving a formal proof of the theorem, we illustrate the proof with a concrete example, from which the general result should be clear. Consider the permutation $\sigma=379218645\in S_{9}$. We look for those $\pi\in\mathcal{P}_{9}$ that satisfy $T_{9}(\pi)=\sigma$. From the definition of the parking process and from the definition of $T_{n}$, we need $\pi_{1}=5$ in order to have $\sigma_{5}=1$, $\pi_{2}=4$ in order to have $\sigma_{4}=2$, $\pi_{3}=1$ in order to have $\sigma_{1}=3$ and $\pi_{4}=8$ in order to have $\sigma_{8}=4$. In order to have $\sigma_{9}=5$, we can either have $\pi_{5}=9$, in which case car number 5 parks in its preferred space 9, or alternatively, $\pi_{5}=8$, in which case car number 5 attempts to park in its preferred space 8 but fails, and then moves on to space 9 and parks. Then we need $\pi_{6}=7$ in order to have $\sigma_{7}=6$. Then similar to the explanation regarding $\pi_{5}$, we need $\pi_{7}$ to be either 1 or 2 in order to have $\sigma_{2}=7$. In order to have $\sigma_{6}=8$, we can have either $\pi_{8}=6$, in which car number 8 parks directly in its preferred space 6, or alternatively $\pi_{8}=5$, in which case car number 8 tries and fails to park in space number 5 and then parks in space number 6, or alternatively, $\pi_{8}=4$, it which case car number 8 tries and fails to park in space number 4 and then also in space number 5, before finally parking in space number 6. Similarly, we need $\pi_{9}$ to be equal to 1,2 or 3 in order to have $\sigma_{9}=3$. Thus, there are $1\times 1\times 1\times 1\times 2\times 1\times 2\times 3\times 3=\prod_{i=1}^{9}\tilde{l}_{9,i}(\sigma)=L_{9}(\sigma)$ different parking functions $\pi\in\mathcal{P}_{9}$ that yield $T_{9}(\pi)=\sigma$.
@@
To give a formal proof for the general case, fix $\sigma\in S_{n}$. In order to have $T_{n}(\pi)=\sigma$, first we need $\pi_{1}=\sigma^{-1}_{1}$. Thus there is just one choice for $\pi_{1}$, and note that $\tilde{l}_{n,1}(\sigma)=1$. Now let $k\in[n-1]$ and assume that we have chosen $\pi_{1},\cdots\pi_{k}$ in such a way that car number $i$ has parked in space $\sigma^{-1}_{i}$, for $i\in[k]$. We now want car number $k+1$ to park in space $\sigma^{-1}_{k+1}$. By construction, this space is vacant at this point, and so are the $\tilde{l}_{n,k+1}(\sigma)-1$ spaces immediately to the left of this space. However the space $\tilde{l}_{n,k+1}$ spaces to the left of this space is not vacant (or possibly this space doesn’t exist–it would be the zeroth space). Thus, by the parking process, car number $k+1$ will park in space $\sigma^{-1}_{k+1}$ if and only if $\pi_{k+1}$ is equal to one of the $\tilde{l}_{n,k+1}(\sigma)$ numbers $\sigma^{-1}_{k+1},\sigma^{-1}_{k+1}-1,\cdots,\sigma^{-1}_{k+1}-\tilde{l}_{n,k+1}(\sigma)+1$. This shows that there are $L_{n}(\sigma)=\prod_{i=1}^{n}\tilde{l}_{n,i}(\sigma)$ different parking functions $\pi$ satisfying $T_{n}(\pi)=\sigma$. $\square$
@@ 3. Proof of Theorem 2
From (3.3) and (3.12), a standard straightforward calculation gives
@@ 4. Proof of Theorem 3
We continue counting in this fashion until we have chosen cars for the $k_{j}-k_{j-1}-1$ spaces between spaces $k_{j-1}$ and $k_{j}$. Recall that $k_{0}$ has been defined to be 0. Now we need for all of the $n-j-\sum_{i=1}^{j}(k_{i}-k_{i-1}-1)=n-k_{j}$ cars that remain from among the first $n-j$ cars to park in the spaces $k_{j}+1,\cdots,n$. In order for this to occur, but for no one of the first $n-j$ cars to park in the $k_{j}$th space, these remaining members of $\{\pi_{i}\}_{l=1}^{n-j}$ must constitute a parking function of length $n-k_{j}$, but shifted forward by $k_{j}$ spaces. This gives $(n-k_{j}+1)^{n-k_{j}-1}$ different choices.
@@
Now we count how many ways we can choice the $\{\pi_{l}\}_{l=n-j+1}^{n}$ in order that $\sigma_{n-j+l}=i_{l},\ l=1,\cdots,j$, or equivalently, in order that car $n-j+l$ park in space $i_{l},\ l=1,\cdots,j$. This calculation uses the same reasoning that was used in the proof of Theorem 1. After the first $n-j$ cars have parked appropriately as above, the spaces $\{k_{l}\}_{l=1}^{j}=\{i_{l}\}_{l=1}^{j}$ are still vacant. From the definition of the parking process and from the definition of $\tilde{l}_{n,n-j+1}(\sigma)$, in order that car $n-j+1$ park in space $i_{1}$, $\pi_{n-j+1}$ must take on one of the values $i_{1}.i_{1}-1,\cdots i_{1}-\tilde{l}_{n,n-j+1}(\sigma)+1$. Then in order that car $n-j+2$ park in space $i_{2}$, $\pi_{n-j+2}$ must take on one of the values $i_{2},i_{2}-1,\cdots i_{2}-\tilde{l}_{n,n-j+2}(\sigma)+1$. Continuing like this, we conclude that there are $\prod_{l=1}^{j}\tilde{l}_{n,n-j+l}(\sigma)$ choices. Using this with (4.1), and noting that there are $(n+1)^{n-1}$ different parking functions of length $n$, we obtain (1.11). $\square$